OpenStudy (anonymous):

What property is shown in the equation? 7ab • 1 = 7ab A. Associative Property of Multiplication B. Commutative Property of Multiplication C. Identity Property of Multiplication D. Reciprocal Property of Multiplication

OpenStudy (bibby):

Associative Property of Multiplication 2*3*4=2*(3*4)=2*12=24 Commutative Property of Multiplication 7*8=8*7 Identity Property of Multiplication 3*1=3 reciprocal 7*\(\frac{1}{7}\)=1

OpenStudy (anonymous):

So I think it ould be Commutative or Associative

OpenStudy (ccswims):

look at identity

OpenStudy (bibby):

this is pattern matching. it's no different than kindergarten when you had to put the red square into the red square hole both of those are wrong

OpenStudy (bibby):

that sounded rude

OpenStudy (anonymous):

oh okay :P ahah its fine i didnt even think it did XD considering our convo before

OpenStudy (anonymous):

im thinkin g c? like @ccswims said

OpenStudy (bibby):

yeah look at it. it states that anything multiplied by 1 remains the same. 7*1=7 -8*1=-8 13728040*1=13728040

OpenStudy (anonymous):

So it is c?

OpenStudy (bibby):

omg i said yeah